# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

### v5

@@ A.3 RQM
Given the challenges in quantifying reference quality through the direct factors mentioned, we instead consider an indirect indication of reference quality using $TNCSI$. Here, $TNCSI$ serves as a quality indicator based on collective user assessments, effectively capturing a statistical summary of numerous researchers’ evaluations of these direct factors for a given paper. Timeliness also matters. An up-to-date literature review ensures that the latest advancements, developments, and perspectives within a field are incorporated. This temporal relevance enhances the accuracy and effectiveness of research outcomes by reflecting the current state of knowledge. By emphasizing the currency of references, we can assess how thoroughly the literature review integrates the most recent research and developments in the field.
To account for both the quality and timeliness of cited references, we propose modifying the Gompertz function to model the reference quality (see in Eq. (3)). The Gompertz function is characterized by its sigmoidal, which indicates a slow growth rate at the start and end of a time period, with more rapid growth in the middle phase. This pattern is often observed in natural phenomena, such as species dynamics [16], tumor growth [118], etc.
The value of $RQM$ is determined by the average reference quality (ARQ), the shift parameter $\beta$, and the median age of references $S_{mp}$. The calculation procedures of $ARQ$ are as follows: The first step is the extraction of the cited reference list. For most publications, their reference lists could be provided by Semantic Scholar API. For a small number of reviews, the reference list provided by Semantic Scholar may contain errors. In this case, although there are powerful computer vision-based algorithms [13, 25] available for extracting the reference list within PDFs, our requirements are relatively simple and can be effectively met by relying on the heuristic algorithms or ChatGPT. More specifically, for literature review with a relatively fixed citation format, we can use the PDFMiner [95] to read text from PDF files and use heuristic rules to match citations. Alternatively, the text can be analyzed using ChatGPT to extract in-text citations. The second step is similar to the approach presented in Sec. 2.4.1, where the ChatGPT and a well-designed prompt (as presented in Fig. 9) are utilized to obtain the topic keyword of the review. Next, we calculate the $TNCSI$ for each reference in the list. To conserve computational resources, we avoid using the ChatGPT to generate keywords for each reference. Instead, the $TNCSI$ of all cited literature is calculated using a sharing topic keyword. Finally, the coverage can be further calculated in Eq. (14):
@@
[…]
@@
Figure 12: 3D visualization of the proposed $RQM$
@@ A.4 RUI
[…]
where $M_{pc}$ denotes the duration in months from the publication of the review to the present, $p=-0.003$, $q=0.001$, $r=0.1267$, $s=0.0129$ are the coefficients obtained by polynomial fitting. Please note that the integral symbol used here is for illustrative purposes only. The strict mathematical definition involves the accumulation of discrete trapezoidal areas.
Finally, the $RUI$ could be obtained by weighted summation of $CDR$ and $RAD$, as shown in Eq. 4. A sculptural visualization of the $RUI$’s contours is crystallized in Fig. 13.
Figure 13: 3D visualization of the proposed $RUI$
A.5 More Examples of Indicator Usage
Tab. XII provides more examples of how the four proposed indicators further assist researchers in refining their review selections. Compared to relying solely on titles, citation counts, and publication dates, the proposed metrics provide more information from various perspectives to assist researchers in review selections.
A.6 Ensuring Fair Use and Avoiding Metric Misinterpretation
@@
A.7 Alignment with the Leiden Manifesto
Principle $TNCSI$ $IEI$ $RQM$ $RUI$
@@ A.7 Alignment with the Leiden Manifesto
TABLE XI: Alignment of Leiden Manifesto Principles with Proposed Metrics: “✓” indicates alignment with the principle, “$\times$” indicates non-alignment, and “N/A” indicates principles not applicable.
Table XI shows the alignment of our proposed four metrics with the specific guidelines of the Leiden Manifesto. The metrics—$TNCSI$, $IEI$, $RQM$, and $RUI$—demonstrate strong adherence to most principles outlined in the manifesto. Specifically, all four metrics align with principles emphasizing the integration of quantitative evaluation with qualitative expert assessment, transparency in data collection and analysis, verification of data and analysis by those evaluated, accounting for field-specific variations, and avoiding misplaced concreteness and false precision.
The metrics do not align with the principle of measuring performance against the research missions of the institution, group, or researcher, as indicated by the “$\times$” in the corresponding row. We would like to point out that this reflects a deliberate trade-off made to balance engineering complexity. While more fine-grained measurements of performance would indeed enhance the objectivity and fairness of the metrics, they would inevitably increase the difficulty of data acquisition. In the future, with advancements in academic search engines and breakthroughs in natural language understanding, this issue is expected to be effectively addressed.
TABLE XII: Comparison of various reviews with the proposed indicators and metrics.
@@ B.3 Review Feature Extraction
Following the discussion in Sec. 3.1, we adopt a cost-effective approach during RAG by limiting the need for the LLM to process the entire text. In simple terms, specific sections of the article are selectively provided to the LLM to generate targeted responses. Alongside the title and abstract, the extracted introduction section and TOC are provided to guide the LLM in identifying whether the authors propose a new taxonomy or include a dedicated section for inclusion and exclusion criteria (indicating adherence to PRISMA guidelines) in the reviewed literature. To assess whether the authors discuss preliminaries, applications, or future challenges, we propose relying exclusively on TOC information, further minimizing costs. For benchmarking, the LLM is instructed to analyze extracted captions (illustrated in Fig. 14) to determine whether the authors have performed quantitative benchmarking of existing methods. All responses generated by the LLM are formatted into a structured 0/1 schema and stored in the RiPAMI database.
@@ Appendix C The Relationship between Review Features and TNCSI
Figure 16: Analysis of six review content features and their association with $TNCSI$: Gene expression heatmap illustrating the independence between six review content features (PRISMA, Application, Benchmark, Preliminary, Taxonomy, and Discussion) and $TNCSI$. Darker shades indicate a match with the feature, corresponding to a value of 1.
Appendix D Code Framework: PyBiblion
This study involved extensive coding and engineering practices during the processes of dataset construction and statistical analysis. To support this work, we developed an open-source code library, PyBiblion, designed for bibliometric and statistical analysis. As its name suggests, PyBiblion is a Python-based library designed for bibliometric and statistical analysis, drawing from the Greek root “biblion”, meaning “book” or “document” to emphasize its focus on scholarly literature.

### v6

@@ A.3 RQM
Given the challenges in quantifying reference quality through the direct factors mentioned, we instead consider an indirect indication of reference quality using $TNCSI$. Here, $TNCSI$ serves as a quality indicator based on collective user assessments, effectively capturing a statistical summary of numerous researchers' evaluations of these direct factors for a given paper. Timeliness also matters. An up-to-date literature review ensures that the latest advancements, developments, and perspectives within a field are incorporated. This temporal relevance enhances the accuracy and effectiveness of research outcomes by reflecting the current state of knowledge. By emphasizing the currency of references, we can assess how thoroughly the literature review integrates the most recent research and developments in the field.
To account for both the quality and timeliness of cited references, we propose modifying the Gompertz function to model the reference quality . The Gompertz function is characterized by its sigmoidal, which indicates a slow growth rate at the start and end of a time period, with more rapid growth in the middle phase. This pattern is often observed in natural phenomena, such as species dynamics [16], tumor growth [121], etc.
The value of $RQM$ is determined by the average reference quality (ARQ), the shift parameter $\beta$, and the median age of references $S_{mp}$. The calculation procedures of $ARQ$ are as follows: The first step is the extraction of the cited reference list. For most publications, their reference lists could be provided by Semantic Scholar API. For a small number of reviews, the reference list provided by Semantic Scholar may contain errors. In this case, although there are powerful computer vision-based algorithms [13, 27] available for extracting the reference list within PDFs, our requirements are relatively simple and can be effectively met by relying on the heuristic algorithms or ChatGPT. More specifically, for literature review with a relatively fixed citation format, we can use the PDFMiner [97] to read text from PDF files and use heuristic rules to match citations. Alternatively, the text can be analyzed using ChatGPT to extract in-text citations. The second step is similar to the calculation of the $TNCSI$, where the ChatGPT and a well-designed prompt (as presented in Fig. 12) are utilized to obtain the topic keyword of the review. Next, we calculate the $TNCSI$ for each reference in the list. To conserve computational resources, we avoid using the ChatGPT to generate keywords for each reference. Instead, the $TNCSI$ of all cited literature is calculated using a sharing topic keyword. Finally, the coverage can be further calculated in Eq. (14):
@@
$\beta_{\text{opt}}=\arg\max_{\beta}\left(\int_{l_{s}}^{r_{s}}\left|RQM^{\prime}(S_{mp};\beta,\overline{ARQ})\right|\,dS_{mp}\right).$ (15)
@@
Figure 15: 3D Visualization of the Proposed RQMRQM
@@ A.4 RUI
$RAD(M_{pc})=\int_{0}^{M_{pc}/12}(px^{3}+qx^{2}+rx+s)\,dx,$ (17)
where $M_{pc}$ denotes the duration in months from the publication of the review to the present, p=0.003p=-0.003, $q=0.001$, $r=0.1267$, $s=0.0129$ are the coefficients obtained by polynomial fitting. Please note that the integral symbol used here is for illustrative purposes only. The strict mathematical definition involves the accumulation of discrete trapezoidal areas.
Finally, the $RUI$ could be obtained by weighted summation of $CDR$ and $RAD$. A sculptural visualization of the $RUI$'s contours is crystallized in Fig. 16.
Figure 16: 3D Visualization of the Proposed RUIRUI
A.5 Ensuring Fair Use and Avoiding Metric Misinterpretation
@@
A.6 Alignment with the Leiden Manifesto
Principle TNCSITNCSI IEIIEI RQMRQM RUIRUI
@@ A.6 Alignment with the Leiden Manifesto
Table 12: Alignment of Leiden Manifesto Principles with Proposed Metrics: ``✓'' indicates alignment with the principle, $\times''$ indicates non-alignment, and ``N/A'' indicates principles not applicable.
Table 12 shows the alignment of our proposed four metrics with the specific guidelines of the Leiden Manifesto. The metrics—$TNCSI$, IEIIEI, RQMRQM, and RUIRUI—demonstrate strong adherence to most principles outlined in the manifesto. Specifically, all four metrics align with principles emphasizing the integration of quantitative evaluation with qualitative expert assessment, transparency in data collection and analysis, verification of data and analysis by those evaluated, accounting for field-specific variations, and avoiding misplaced concreteness and false precision.
The metrics do not align with the principle of measuring performance against the research missions of the institution, group, or researcher, as indicated by the $\times''$ in the corresponding row. We would like to point out that this reflects a deliberate trade-off made to balance engineering complexity. While more fine-grained measurements of performance would indeed enhance the objectivity and fairness of the metrics, they would inevitably increase the difficulty of data acquisition. In the future, with advancements in academic search engines and breakthroughs in natural language understanding, this issue is expected to be effectively addressed.
A.7 Illustrative Applications of the Indicators
Table. 13, Tab. 14 provide more examples of how the four proposed indicators further assist researchers in refining their review selections. Compared to relying solely on titles, citation counts, and publication dates, the proposed metrics provide more information from various perspectives to assist researchers in review selections.
Table 13: Comparison of Various Reviews with the Proposed Indicators and Metrics (Part 1).
[…]
Table 14: Comparison of Various Reviews with the Proposed Indicators and Metrics (Part 2, continued).
@@ B.3 Review Feature Extraction
We adopt a cost-effective approach during RAG by limiting the need for the LLM to process the entire text. In simple terms, specific sections of the article are selectively provided to the LLM to generate targeted responses. Alongside the title and abstract, the extracted introduction section and TOC are provided to guide the LLM in identifying whether the authors propose a new taxonomy or include a dedicated section for inclusion and exclusion criteria (indicating adherence to PRISMA guidelines) in the reviewed literature. To assess whether the authors discuss preliminaries, applications, or future challenges, we propose relying exclusively on TOC information, further minimizing costs. For benchmarking, the LLM is instructed to analyze extracted captions (illustrated in Fig. 17) to determine whether the authors have performed quantitative benchmarking of existing methods. In addition, the LLM is guided to judge whether the abstract follows structured-abstract standards [12]. All responses generated by the LLM are formatted into a structured 0/1 schema and stored in the RiPAMI database.
@@
B.4 Position Statistics of Section Titles
To analyze the structural characteristics of reviews, we compute the normalized positions of section titles that match a predefined set of keywords (e.g., introduction, conclusion). The procedure follows three steps.
First, the keyword list is standardized by removing empty entries, converting to lowercase, and eliminating duplicates while preserving the original order. Second, for each paper, all section titles are retrieved, and the total number of titles is used to normalize their positions within the document. Specifically, the index of each title is divided by (n1)(n-1), where $n$ denotes the number of titles in the paper, yielding a value between 0 (beginning of the document) and $1$ (end of the document). Finally, each title is checked against the keyword list in a case-insensitive substring matching manner, allowing one title to be counted under multiple keywords if applicable.
This method produces, for each keyword, a distribution of normalized positions across the entire corpus, which forms the basis for our structural analysis.
@@ Appendix C The Relationship between Review Features and TNCSI
Figure 19: Analysis of Six Review Content Features and Their Association with TNCSITNCSI: Gene expression heatmap illustrating the independence between six review content features (PRISMA, Application, Benchmark, Preliminary, Taxonomy, and Discussion) and $TNCSI$. Darker shades indicate a match with the feature, corresponding to a value of 1.
Appendix D Complete List of Subfields in CV, NLP, and MISC
For transparency and reproducibility, we provide a complete catalogue of subfields across the three major domains, namely CV, NLP, and MISC, as summarized in Table 15.
Table 15: Comprehensive Subfield Catalogue of CV, NLP, and MISC Domains
Domain Subfields
CV action detection, action recognition, activity detection, activity recognition, anomaly detection, boundary detection, cnn, computer vision, depth estimation, edge detection, emotion recognition, face detection, face recognition, facial recognition, gesture analysis, gesture recognition, hand gesture recognition, handwriting recognition, human activity recognition, human detection, human pose estimation, image captioning, image classification, image clustering, image compression, image editing, image enhancement, image generation, image inpainting, image matching, image quality assessment, image recognition, image reconstruction, image restoration, image retrieval, image segmentation, image-based localization, instance segmentation, medical image analysis, medical image segmentation, object detection, object tracking, optical character recognition, person re-identification, point cloud, saliency detection, salient object detection, scene segmentation, scene understanding, super-resolution, superpixels, video object segmentation, video processing, video summarization, video understanding, visual question answering, visual tracking
NLP dialogue modeling, dialogue systems, document analysis, document analysis and recognition, document clustering, document layout analysis, document retrieval, language modeling, language modelling, machine translation, named entity disambiguation, named entity recognition, natural language processing, question answering, relation extraction, sentiment analysis, sentiment classification, text classification, text clustering, text generation, text mining, text summarization, text-to-image generation, text-to-speech conversion, text-to-speech synthesis
MISC adversarial attack, audio classification, biometric authentication, biometric identification, contrastive learning, data mining, data visualization, diffusion model, domain adaptation, graph mining, knowledge graph, knowledge representation, machine learning interpretability, meta-learning, metric learning, multi-label classification, pattern matching, pattern recognition, pre-training, pretraining, prompt learning, recommendation systems, recommender systems, remote sensing, representation learning, self-supervised learning, semantic segmentation, signature verification, speech emotion recognition, speech enhancement, speech recognition, speech synthesis, speech-to-text conversion, time series analysis, time series forecasting, topic detection, topic modeling, transfer learning, unsupervised learning, vision language model, word embeddings, zero-shot learning
Appendix E Code Framework: PyBiblion
This study involved extensive coding and engineering practices during the processes of dataset construction and statistical analysis. To support this work, we developed an open-source code library, PyBiblion, designed for bibliometric and statistical analysis. As its name suggests, PyBiblion is a Python-based library designed for bibliometric and statistical analysis, drawing from the Greek root ``biblion'', meaning ``book'' or ``document'' to emphasize its focus on scholarly literature.